OpenStudy (anonymous):

What is the slope of a line perpendicular to the line 12x - 4y = 7? A. -7/4 B. 7/4 C. -1/3 D. 1/3

OpenStudy (shamil98):

The slope of line perpendicular is a negative reciprocal. First, put your equation into y =mx + b form. To get your slope ( the m variable in this equation).

OpenStudy (anonymous):

3x-7= y thats your stand mx+c = y form. Slope of the perpedicular line will be -1/3

OpenStudy (anonymous):

3x- 7/4*

OpenStudy (anonymous):

Could you help on another as well?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Write an equation of a line in slope-intercept form that is parallel to x + 4y = 6 and passes through the point (-8, 5) A. y=−14x+3 B. y=−14x+6.75 C. y = x - 13 D. y = x + 13

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

for paralell you slope doesnt change.. cause _ its paralell

OpenStudy (anonymous):

-1/4 will be the slope

OpenStudy (anonymous):

so it would be B?

OpenStudy (anonymous):

@utkrij